solve the following using factor theorem x^3+6=2x^2+5x
Answer #1

using factors theorem x²+6= 2 x 750 x3 2x²5x+6=0 => f(x)= x ² 2x²5x+6=0. Take azl fedz 1-2-5+6 27-70 .. n=1 is a not of f(x)